OpenStudy (anonymous):

4< x < 1/2 x could equal to?

OpenStudy (anonymous):

Answers 40% None of the above 45% 50%

OpenStudy (anonymous):

what? why would the answer be written as a percent?

OpenStudy (anonymous):

first of all before doing any computation, let's think: if \[x<\frac{1}{2}x\]is it clear that x has to be negative?

OpenStudy (anonymous):

because the only way for a number to be less than half of itself, it certainly cannot be positive! it does however work with negative numbers. any negative number is less that half of itself. for example -4 is less than -2

OpenStudy (anonymous):

but you also have that x is greater than 4. so x must be positive. therefore this inequality makes no sense, i.e. there is no solution

OpenStudy (anonymous):

unless i am a total idiot and this actually just says \[4<x<\frac{1}{2}\] in which case it is ridiculous because there is no way that \[4<\frac{1}{2}\]

OpenStudy (anonymous):

true